if cos x = -and sin x > 0, find the exact value of sin(2x) VO and simplify
Answer #1

solu no cosse + Sim x = T-cos?re= sitten * sinus id = 15 - sinu = FS XJ6 = 130 TEXTO no Sin 29= 2sina.cosx = 2* 5) Sibase = 1